News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 5 October 2001 3.30pm EDT

During last week the observing schedule was interrupted at 2.40am on Oct 2 by a solar radiation event that stopped the daily load and activated the science instrument safing sequence (SCS 107). All actions were nominal and assessment of the dump data indicated that the trigger was due to count rates in all of the EPHIN channels exceeding their pre-defined threshold levels. The loads were restarted at 8am on October 3 after radiation levels returned to acceptable levels. The observations of UV Ceti, IRAS 05189-2524 and the Puppis bright E knot were impacted by the event and will be rescheduled in a future load.

A number of real-time procedures were completed during the week relating to the solar interruption and re-starting of science. These included starting an ACIS CTI measurement, re-enabling SCS 107 and executing a momentum unloading.

The rescheduled portion of this week and next week's schedule is shown below.

------------------------------------------
Cal 83                  HRC-S/LETG   Oct 3
Radiation Belts                      Oct 4
Abell 496               ACIS-S
M31 Nucleus             ACIS-S
M31 Nucleus             ACIS-I       Oct 5
A2556                   ACIS-S
SAX J1753.5-2349        ACIS-S
XTE1650-500             ACIS-S/HETG
NGC6093                 ACIS-S       Oct 6
Radiation Belts
RCW103                  ACIS-S       Oct 7
PKS 2135-14             ACIS-S
MS 2053.7-0449          ACIS-I

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports